### v1

# DataPlatter: Boosting Robotic Manipulation Generalization with Minimal Costly Data

Abstract
The growing adoption of Vision-Language-Action (VLA) models in embodied AI intensifies the demand for diverse manipulation demonstrations. However, high costs associated with data collection often result in insufficient data coverage across all scenarios, which limits the performance of the models. It is observed that the spatial reasoning phase (SRP) in large workspace dominates the failure cases. Fortunately, this data can be collected with low cost, underscoring the potential of leveraging inexpensive data to improve model performance. In this paper, we introduce the DataPlatter method, a framework that decouples training trajectories into distinct task stages and leverages abundant easily collectible SRP data to enhance VLA model’s generalization. Through analysis we demonstrate that sub-task-specific training with additional SRP data with proper proportion can act as a performance catalyst for robot manipulation, maximizing the utilization of costly physical interaction phase (PIP) data. Experiments show that through introducing large proportion of cost-effective SRP trajectories into a limited set of PIP data, we can achieve a maximum improvement of 41% on success rate in zero-shot scenes, while with the ability to transfer manipulation skill to novel targets.
footnotetext: 1 Meituan Inc.footnotetext: 2 Institute of Computing Technology, Chinese Academy of Sciences.footnotetext: Corresponding authors.
1 Introduction
As the understanding and reasoning abilities of Multi-modal Large Language Models (MLLMs) advance rapidly, their application in real-world interactions, i.e. Embodied Artificial Intelligence (EAI), has become a focal point of research [27, 4, 14], and the method utilizing Vision-Language-Action (VLA) models is a common choice [16, 53, 47, 5]. Similar to MLLMs, training the spatial understanding and physical interaction reasoning abilities of VLA requires a large quantity of demonstration trajectories across a variety of tasks. Although much effort at high cost has been dedicated to collecting robot demonstrations, both in simulation [10, 30, 11] and the real world [3, 35, 41], generalizing agent-specific trajectories to a novel agent configuration remains a critical challenge. As a result, the training data available for a specified agent remains limited, which is far from sufficient to encompass the diverse real-world scenarios, thereby constraining the improvement of the VLA models’ capabilities.
Figure 1: Motivation for Our Work: (a) Demonstration trajectories used in embodied AI training are typically collected via teleoperation, which is both time-consuming and expensive. (b) However, most task trajectories can be segmented into a spatial reasoning phase (SRP) and a physical interaction phase (PIP), each with distinct focus and learning difficulty. (c) The SRP data can be automatically collected using simple algorithms at a high speed.
To address this issue and enhance data utilization efficiency, researchers are focusing on exploring cross-agent training [47, 6, 35, 44, 23], spatial cognition enhancement [25, 12, 51] and task logical extraction [38] through chain-of-thoughts. Notably, recent studies [41, 24] have demonstrated a scaling law governing the relationship between the spatial volume of operation workspace, the quantity of training data and the generalization performance of VLA models. All these approaches share a common premise: understanding the compositional nature of embodied tasks.
Through analyses we reveal that most tasks process can generally be divided into two stages: the Spatial Reasoning Phase (SRP) and the Physical Interaction Phase (PIP), as shown in Fig. 1. The former stage is target-agnostic, as the agent explores extensive workspace without any close interaction with the targets, such as approaching the target before operation, making data collection relatively straightforward. In contrast, during the later stage, precise actions governed by physical laws should be applied to the target with the foresight of object reaction, which is extremely labor-intensive, either for human or algorithmic experts. This motivates our core question: can inexpensive SRP data amplify the value of scarce PIP data thus reduce the effort required for data collection?
Our key insight stems from two critical observations: (1) The spatial understanding ability required in SRP exhibits higher environmental variability compared to PIP, since the manipulation stage for a specified target is relatively fixed with little correlation with the surrounding scene; (2) Neural networks demonstrate distinct attention patterns during different task stages, such as the focus on target’s location and spatial occupancy to avoid collision in SRP while shifted to the target proportion in PIP. These findings suggest that sub-task-specific training strategies could better align with the model’s learning characteristics, utilizing varying proportions of these data segments across the sub-tasks.
Both Tan et al. [41] and our experiments (see Tab. 3) have demonstrated that a smaller workspace can significantly improve the success rate of operating tasks. This indicates that decoupling operation stages with different centers of attention can improve generalization performance. Furthermore, this variation in sub-task difficulty can cause the model to overfit on the simpler, small-workspace stage while underfit on the large-workspace stage, which necessitates different data volumes for each stage. In this paper, we propose the DataPlatter method, which decouples training data across different operation stages, constructs a implicit sub-task specific training procedure, and leverages a large amount of easily collectible SRP data to train this stage, to improve the performance of VLA models. With our method, a substantial amount of labor-intensive teleoperation time traditionally required to collect complex manipulation trajectories, e.g. the 17 months needed for 130k episodes in RT-1 [3], can be significantly reduced. Instead, program-driven automatic collection can be employed to acquire a large volume of low-interaction trajectories in extensive workspaces. This approach not only reduces manual efforts but also greatly enhances the potential to leverage larger datasets to improve model capabilities.
The contribution of this paper is as follows:
  • We introduced the DataPlatter methods, which utilizes additional cost-effective SRP trajectories to improve the model’s generalization performance in zero-shot scenes.
  • We prove that SRP data can act as catalyst to maximize the contribution of expensive manipulation dataset in VLA model training.
  • Decoupling task stages in a dataset to build implicit sub-task specific training processes can offer a flexible approach to enhancing model performance at the sub-task level.
  • Experiments demonstrate that our method increases the task success rate by 41% in zero-shot scenarios and can effectively transfers model skills to novel target objects.
2 Related Work
Multi-modal VLA models
Unlike previous studies [34, 31, 48] that employed models of limited size and do not heavily reliant on large volumes of training data, recent research efforts such as RoboMM [47], RoboFlamingo [22] and $\pi_{0}$ [2] have leveraged MLLMs to achieve a generalist performance across multiple long-horizontal tasks through Imitation Learning (IL). Consequently, these approaches necessitate a substantial amount of data, imposing significant challenges in data collection. Numerous studies have invested considerable efforts in training with multiple datasets [35, 16, 6, 44]. However, generalization across different tasks, embodiments, and datasets remains a significant challenge, necessitating further fine-tuning on specific datasets during evaluation. Another line of research utilizes pre-training with easily obtainable data formats [5, 25, 53] to capture knowledge of the world, but still requires a large volume of action data to perform specified tasks effectively. Furthermore, diffusion-based methods [49, 15, 21, 37, 28], as well as Vector-Quantization (VQ) methods [33, 40, 18], demand substantial amounts of action trajectories to adequately encapsulate high-dimensional probability distributions and codebooks. This paper proposes a data mixture method that reduces the reliance on costly manipulation trajectories, offering a partial solution to the aforementioned challenges.
Robot manipulation datasets
The EAI community have released a number of large-scale datasets collected in both simulation [10, 30, 50, 11, 19] and real world [35, 46, 3, 43, 9]. However, most datasets are collected through teleoperation and manual labeling, which is an extremely time-consuming process. Furthermore, the configurations of the embodiments, tasks and scenes in these datasets are different, posing challenges in reproducing performance in local experiments, particularly for datasets collected in real-world settings. On the other hand, datasets collected through algorithm-driven methods, which are primarily gathered in simulators using fixed task templates [50, 29] or Reinforcement Learning (RL) with task disassembly [45, 11, 42], are suffering from a lack of task diversity and often involve simplified physically simulations that are impractical for real-world deployment. With our method, models can be trained with a large proportion of easily collectible trajectories, which can be automatically collected through much simpler process, reducing the models’ need for expensive interaction data.
Generalizing model capability
Currently, most EAI models are limited to executing tasks they have explicitly encountered during training. For instance, even if a model be trained to pick up bottles, it cannot generalize this to pick up a cola can. Although this problem have already been studied through methods ranging from early domain randomization [13], meta-learning [8] and data augmentation[17] to recent advancements in world model building [5, 27] and spatial reasoning [25, 12], the generalization performance on out-of-distribution (OOD) novel targets still shows limited improvement. [5, 53] try to transfer the world knowledge from large models trained with Internet-scale data to robot action reasoning, but the the manipulation experience of OOD targets from “practicing” can not be efficiently acquired from “reading”, while [27, 38, 38] are trying to directly use the general ability to guide the agent’s action logic. [25, 12, 51] are working improving the action performance through understanding the spatial information in the workspace. Zhu et al. [52] transfer the target knowledge to similar objects through text-image pairs, but still needs auxiliary information to get a better performance during inference. In this paper we propose an end-to-end training method, which can improve the generalization performance on OOD targets by a large margin.
3 Method
As illustrated in Fig. 2, this paper introduces the DataPlatter method, which segments robot manipulation trajectories into spatial reasoning and physical interaction phases according to the extent of the agent’s interaction with objects in the environment. By employing a mixture of two-stage data in appropriate proportions, we aim to achieve a generalization performance comparable to using complete data for model training. This approach effectively reduces the reliance on expensive PIP data.
Figure 2: Overview of our method. We divide each training trajectory into two stages: the spatial reasoning phase (SRP), which does not require close interaction, and the physical interaction phase (PIP), during which the agent directly operates the target. $N_{1}$ SRP and $N_{2}$ PIP trajectories are sampled to form a new dataset (left). This resampled dataset is used to train the VLA model in an IL manner (middle). During the inference process, the VLA model takes the agent’s observation and language instruction as input and predicts the next step action to guide the agent in accomplishing the task (right).
3.1 Problem Formulation
In this paper we focus on the VLA models that utilize behavior cloning, which is a category of IL methods. Consider a robot manipulation trajectory dataset $\mathcal{D}^{F}=\left\{\tau_{i}^{F}\right\}_{i=1}^{N}$, where each full-stage trajectory $\tau_{i}^{F}=\left\{l^{i},o_{1}^{i},a_{1}^{i},o_{2}^{i},\cdots,a_{T-1}^{i},o_{T}^{i}\right\}$ consist of the task’s language instruction $l$, the agent’s observation $o_{t}^{i}$ at each time step $t$, and the action $a_{t}^{i}$ taken by the agent. The VLA model $\Psi_{\theta}$ with parameter $\theta$ takes as input the task instruction and a segment of observation history $\mathcal{O}_{t,L}^{i}=\left\{o_{t-L+1},\cdots,o_{t}\right\}$ of length $L$, and predicts the action chunk $\mathcal{A}_{t,L,H}^{i}=\left\{a_{t-L+1},\cdots,a_{t+H}\right\}$ that the agent should execute to accomplish the task in the past $L$ and next $H$ time steps, i.e.
$\hat{\mathcal{A}}_{t,L,H}^{i}$ $=\Psi_{\theta}\left(\mathcal{O}_{t,L}^{i},l^{i}\right)$ (1)
$=Dec\left(LLM\left(Enc_{v}\left(\mathcal{O}_{t,L}^{i}\right),Enc_{l}\left(l^{i}\right)\right)\right),$
where $Enc_{v}$ and $Enc_{l}$ are vision and language encoders, $LLM$ is the pretrained large language model, and $Dec$ donates the action decoder. Typically, vision encoders like CLIP [36] are pretrained using image-text pairs to provide aligned visual-textural semantics, facilitating seamless integration with LLMs, and are generally kept frozen during training of the VLA model. The LLM such as GPT [1] or LLaMA [32] serves as the core of the model due to its robust general reasoning capabilities and typically employs adapters [20, 26] for integrating multi-modal input tokens. Action decoders usually consist of several lightweight neural layers that interpret the action token chunks output by the LLM and transform them into physically meaningful actions, e.g. 6-DoF poses of end effectors.
The objective of model optimization is to minimize the discrepancy between predicted and demonstration action sequences, i.e.
$\theta^{*}=\arg\min_{\theta}Err\left(\hat{\mathcal{A}}_{t,L,H}^{i},\mathcal{A}_{t,L,H}^{i}\right),$ (2)
where $Err\left(\cdot\right)$ is a function that measures the divergence between the demonstrated and predicted actions, and has different representations for different action forms.
3.2 VLA Training with DataPlatter
To utilize the trajectories of different sub-tasks in the dataset, we first segment the given full-stage trajectory $\tau_{i}\in\mathcal{D}$ into SRP and PIP based on the distance between the end-effector $G$ and the target object $T$, as well as the visibility of the target in the wrist camera $C_{w}$. It is assumed that there is one wrist camera and one static camera in the scene, which is a common configuration in most datasets. More formally, for a target object $T$ at position $p_{T}$, an end-effector $G$ at position $p_{G}$ and a wrist camera $C_{w}$ at pose $P_{C}=\left(p_{C},R_{C}\right)$ defined under OpenCV frame, where $p_{T},p_{G},p_{C}\in\mathbb{R}^{3}$ and $R_{C}\in SO(3)$, the PIP begins if
$\left\{\begin{matrix}\left\|p_{G}-p_{T}\right\|\leq d_{th},\\ \frac{v_{z}^{T}\cdot\left(p_{T}-p_{C}\right)}{\left\|p_{T}-p_{C}\right\|}>\arccos\frac{\alpha_{fov}}{2},\end{matrix}\right.$ (3)
where $v_{z}\in\mathbb{R}^{3}$ s.t. $\left\|v_{z}\right\|=1$ is the direction vector of $z$-axis of frame $P_{C}$, $d_{th}$ is the distance thresholds and $\alpha_{fov}$ is the field of view of $C_{w}$. The PIP stops once the interaction-rich manipulation stage is accomplished, e.g. after grasping the target in pick-and-place task or triggering the button in switch-operation task. Apart from the PIP, the rest of the trajectory is designated as the SRP. Following such procedure, the trajectory can be divided into several segments $\tau_{i}^{F}=\left\{\tau_{i,1}^{SRP},\tau_{i,1}^{PIP},\tau_{i,2}^{SRP},\cdots\right\}$. Correspondingly, the dataset can be divided into two sub-datasets: $\mathcal{D}^{F}=\mathcal{D}^{SRP}\cup\mathcal{D}^{PIP}$, where $\mathcal{D}^{SRP}=\left\{\tau_{i,j}^{SRP}\right\}$ contains all of the SRP segments in the trajectories and $\mathcal{D}^{PIP}=\left\{\tau_{i,j}^{PIP}\right\}$ contains the manipulation segments. Note that we are aiming to train the VLA model with a larger quantity of easily-collectible SRP data than expensive PIP data, so in practice independently-collected SRP dataset $\mathcal{D}_{ind}^{SRP}$ can be included in training.
Before the training phase of the VLA model, we sample $N_{1}$ and $N_{2}$ segments in $\mathcal{D}$ and $\mathcal{D}_{ind}^{SRP}$ respectively, and construct a new dataset $\mathcal{D}^{Mix}$ to train the model, which in this paper we call DataPlatter, i.e.
$\mathcal{D}^{Mix}=\left\{\tau_{i}^{F}\sim\mathcal{D}^{F}\right\}_{i=1}^{N_{1}}\cup\left\{\tau_{i}^{SRP}\sim\mathcal{D}_{ind}^{SRP}\right\}_{i=1}^{N_{2}}.$ (4)
In practice, to achieve the best model capability, generally the whole full-stage trajectory dataset $\mathcal{D}^{F}$ is used, i.e. $N_{1}=\left|\mathcal{D}^{F}\right|$, and select a proper $N_{2}$ to improve the generalization performance on novel scenes. Through this method, a implicit sub-target specific training with sub-task datasets $\mathcal{D}^{PIP}$ and $\mathcal{D}^{SRP}\cup\mathcal{D}_{ind}^{SRP}$ is constructed, providing a flexible way to control the performance of each sub-task. By varying the proportion of data between the two sub-datasets, a tendency in the task success rate relative to the amount of SRP data can be observed in Sec. 4, from which a principle for conserving PIP data while maintaining the VLA model’s performance can be concluded.
4 Experiments
In this section, we investigate how the total number of training trajectories and the proportion of SRP data impact task success rates. We aim to identify the optimal strategy for leveraging easily accessible SRP data to enhance the generalization performance of the VLA model.
4.1 Environment Setup
VLA model
Unless otherwise stated, in this paper the RoboMM [47] is used as our baseline, which is a multi-modal VLA model that utilizes UVFormer [25] to help with spatial perception through RGB image with camera parameters in a low-cost manner. During training we feed language instruction and RGB images from a static camera and a wrist camera, together their intrinsic and extrinsic parameters, into the model, and use the depth images with action chunks as supervision.
Training data
In the simulation environment of IsaacSim, we generated a dataset for object-picking tasks involving target objects of various categories and geometrical shapes. For SRP-only trajectories, to provide a implementable pipeline in real-world robots, we did not directly read the object information from simulation, instead we applied a detection-sampling method provided in Fig. 1(c). Details are shown in Sec. A.1. During trajectory generation, we observe that the SRP-only trajectories are generated 2.5×2.5$\times$2.5 × faster than those using full-stage data, while the length of the full-stage data is only 1.4×1.4$\times$1.4 × that of the SRP data. In real-world data collection this discrepancy can only be even larger. Other datasets we used in the experiments are divided using method provided in Eq. 3 with $d_{th}=0.2m$ and $\alpha_{fov}=\frac{\pi}{3}$, details are shown in Sec. A.2.
Evaluation
We evaluated our models in the aforementioned simulation environment. A trail is considered succeeded if the agent successfully picked up the instruction-specified target object under the actions generated by the VLA model. During evaluation the scenes are divided into test and zero-shot configurations. The test scenes are configured in ways that have been encountered during training, while the zero-shot scenes are initialized randomly to test the generalization performance. Note that depth images are not utilized in the model inference, only language instruction, RGB observation and camera parameters are fed into the model.
Training
Our models are trained on servers equipped with 8 Nvidia A100 GPUs, each with 80GB of CUDA memory. The SRP segments are generally longer than the PIP segments, and the dataset $\mathcal{D^{Mix}}$ contains several times more SRP trajectories compared to PIP trajectories. During training we form the mixed dataset $\mathcal{D}^{Mix}$ with a varying of proportion of independent SRP segments, i.e.
$p_{SRP}=\frac{N_{2}}{N_{1}+N_{2}}.$ (5)
To prevent the SRP features from dominating the model’s understanding of operations, during training, the PIP trajectories $\tau^{PIP}\subset\tau^{F}$ are duplicated $\left\lfloor\frac{N_{2}}{N_{1}}\right\rfloor$ times. The benefits of this approach are discussed in Sec. 4.3.3. The checkpoint with the best performance in zero-shot environments within the first 10 epochs is used for evaluation.
4.2 Experiment Results
Figure 3: Impact of training trajectories quantity per task on mean success rate among all tasks in zero-shot scenes. The reference model is trained using with a configuration of $N_{1}=100,N_{2}=0$ to serve as the baseline performance. The horizontal axis represents the number of training trajectories for a single task, i.e. $N_{1}+N_{2}$.
4.2.1 Generalization with SRP Data
Firstly, we aim to determine whether increasing the amount of SRP data can improve the model’s generalization performance. We randomly select the PIP trajectories of $M$ target objects as our task, and create a multi-task dataset to train the RoboMM models. For each task, we use a fixed number of $N_{1}=100$ full-stage trajectories $\tau_{i}^{F}\in\mathcal{D}^{F}$, alongside a variable number of $N_{2}$ SRP trajectories $\tau_{i}^{SRP}\in\mathcal{D}_{ind}^{SRP}$. In this experiment, we set $M\in\left\{5,20\right\}$ and $N_{2}\in\left\{0,100,200,300\right\}$. The results are presented in Fig. 3.
It can be observed that incorporating additional automatically collected SRP data during training significantly elevates the model’s success rate, achieving a maximum margin of 41% over the reference model. Meanwhile, incorporating SRP data can achieve an equivalent performance to manually collected full-stage data (marked with gray arrow) at substantially lower cost. However, the performance bottleneck occurs earlier than that of using full-stage data, when $N_{2}>2N_{1}$, i.e. in this experiment $N_{2}>200$, increasing the number of SRP trajectories $N_{2}$ while keeping $N_{1}$ constant yields no significant improvement in the model’s performance. This suggests that incomplete trajectories cannot be added indefinitely, as insufficient operational data may hinder the model from learning effective action logic, and the proportion of data mixture is discussed in Sec. 4.3.2. However, at this point, the generalization performance of a model trained with $N_{1}=300$ full-stage trajectories across 5 tasks can still improve by 7% ($68\%\rightarrow 75\%$) with an additional $N_{2}=300$ SRP trajectories. This performance is close to the bottleneck of 80% achievable with more full-stage data.
4.2.2 Universality of SRP Data
To verify that adding SRP data can improve the model performance across different models and datasets, we trained and tested RoboMM [47] and RoboFlamingo [22] on both our dataset and CALVIN dataset [30] using the aforementioned approach. The results are shown in Tab. 1, in which the “w/ SRP” refers to models trained with SRP data at $p_{SRP}=66\%$, while “w/o SRP” donates models trained with the same $N_{1}$ as the former but $N_{2}=0$, and “FS” donates the model trained with totally full-stage trajectories, to serve as the upper bound of model performance.
It can be observed that in all of the settings, adding additional SRP segments data can significantly improve the generalization performance in different models and different tasks. The improvement amplitude compared to the baseline model (marked w/o SRP) on our dataset is significantly higher than on CALVIN. This is because the tasks in CALVIN are simpler, while the trajectories approaching the target in our dataset are more complex. In our dataset, agents approach the target from random orientations rather than a relatively fixed pose, as in CALVIN, resulting in a larger search space for the SRP policy. These findings underscore the advantage of our method in handling tasks with expansive workspaces.
Dataset Model w/o SRP w/ SRP FS
Ours RoboMM 0.28 0.58 0.68
Ours RoboFlamingo 0.13 0.28 0.44
CALVIN RoboMM 0.80 0.88 0.93
CALVIN RoboFlamingo 0.74 0.78 0.81
Table 1: Model performance with additional SRP segments on multiple models and datasets.
4.2.3 Generalization on Novel Target
Additionally, for tasks requiring a deep understanding of object geometry and physical laws, such as picking up objects with totally different geometries, we wonder whether the VLA models can transfer generalized knowledge to out-of-distribution (OOD) target objects. To verify this, we introduce new tasks featuring novel target objects and only SRP data during training, within datasets containing other targets with different geometries and totally full-stage trajectories, and see that if the novel task can be successfully executed. The results are shown in Tab. 2.
With only SRP data, the models can successfully pick up the novel target even if they have not seen examples on how to do this, especially for the 5 tasks with similar geometry (see Fig. 8), only through the experience on the other targets. Meanwhile, without SRP segments, the models are just wandering in the workspace without knowing what to do. This result indicates that the similarity between the SRP segments of the novel targets and the others acts as a bridge, enabling the models expand the skill of the entire task to the novel target, without requiring any additional auxiliary information. This finding significantly broadens the scope for future research on task generalization performance. However, performance on novel targets remains relatively low, indicating that further research is needed in this area.
# Tasks Config. Seen Targets Novel Targets
4 Seen &1 Novel None 0.67 0.05
SRP 0.65 0.40
SRP + PIP 0.64 0.65
10 Seen &10 Novel None 0.62 0.03
SRP 0.61 0.20
SRP + PIP 0.74 0.65
Table 2: Generalization performance on OOD target objects. Performance of in-distribution targets are also presented as a control. Data configuration of novel target: “None” - No training data on novel target objects; “Only SRP” - Novel target objects have only SRP segments in the dataset; “SRP + PIP” - Novel objects are trained using the same full-stage data as other tasks.
4.2.4 Failure Analysis
During the experiments, we observed that the most common failure cases in tasks utilizing additional SRP data was the agent attempting to operate the target from an unreasonable pose, making it unable to pick up the target, as shown in Fig. 4. Besides, another major reason is that the target frequently slipped from the agent’s fingers due to a pick action with too shallow bite. However, in the models trained with full-stage trajectories, misidentifying the correct target was the most frequent failure (42.6%) aside from the aforementioned, while it was rarely observed in tasks utilizing additional SRP data. This phenomenon demonstrates that we have successfully developed a training process tailored to implicit sub-tasks, achieving specialization through focused training. This result indicates that the performance of a specific task stage can be enhanced by incorporating independent data relevant to itself without adversely affecting other stages.
Figure 4: Most common failure reasons for tasks trained with additional SRP data. (Left) Unreasonable manipulation pose, in this example the gripper are supposed to manipulate the target from the thinnest direction. (Right) Too shallow bite, leading to unstable grasp result.
4.3 Ablation Studies
4.3.1 Stage Decoupling Necessity
To assess the impact of the large-workspace SRP stage on model performance, we trained two models: one including the data of SRP segment and one without it. Both models were evaluated in environments initialized with configurations either before or after approaching the target. In the testing environments where the agent is initialized near the target, the end effector is randomly positioned within a maximum range of $0.15m$ from the target object, as illustrated in Fig. 1(c). In the alternative setup, the agent starts in a random pose within the workspace. The results are shown in Tab. 3.
Train w/ SRP Evaluation Test Scenes Zero-Shot
Random 0.53 0.40
Near Target 0.80 0.68
$\times$ Random 0.08 0.06
$\times$ Near Target 0.73 0.66
Table 3: Influence of the SRP stage on the model performance.
In both models, the evaluation revealed that the SRP stage significantly reduced performance. Although the SRP stage provides the necessary capability for the model to locate the target, its difficulty is much higher than that of the PIP stage. This increased difficulty stems from the vast exploration space, which introduces diverse situations. Consequently, it is advisable to learn the two stages in a decoupled manner, i.e. train the two sub-tasks with different amount of data at varying speed, which is the the way our method employs.
To verify the impact of the wide-range workspace on task difficulty during the SRP stage, we trained two sets of models: one using full-stage trajectories and the other using only PIP segments. The results, as shown in Fig. 5, indicate that the performance bottleneck, i.e. the data volume required for the model to converge to a relatively stable success rate, occurs much earlier in the simpler manipulation-only tasks compared to the full-stage tasks. Together with the scaling law related to the target variety shown in Fig. 6, we observe that the difficulties of different sub-tasks arise from distinct aspects: the space-related SRP stage requires demonstrations that cover the entire workspace, whereas the geometry-related PIP stage necessitates a variety of target shapes to derive an effective manipulation strategy. These findings highlight the necessity of preparing decoupled data for the various stages within a single task category.
Figure 5: The scaling relationship between task performance and the data volume for models trained with and without the data of SRP segments.
Figure 6: The scaling relationship between task performance and the number of target objects with various geometry shapes.
4.3.2 Scaling with SRP Proportion
To determine the proportion of inexpensive SRP trajectories $p_{SRP}$ that can be incorporated into the model without significantly impacting performance, we trained several models on 5 tasks, each with a total of $N_{1}+N_{2}=300$ trajectories. The results are presented in Fig. 7. It can be observed that there is a logarithmic relationship between the the proportion of SRP data and the model success rate, expressed as $SR=k\ln\left(1-p_{SRP}\right)+b$, and when $P_{SRP}>66\%$, the rate of decline in mean success rate increases rapidly. This decline is due to the increasing reliance on trajectories that lack a manipulation phase, which undermines the model’s object manipulation skills, causing the agent to merely wander around the target. In particular, when $\frac{dRS}{dp_{SRP}}=\frac{k}{p_{SRP}-1}=-1$, i.e. $p_{SRP}=1-k\approx 80\%$ in our experiments, the rate of performance decline exceeds that of increasing the proportion of independent SRP data. This result suggests us that for a existing dataset $\mathcal{D}^{F}$ containing $N_{1}$ full-stage trajectories, we can add at most $N_{2}=\left|\mathcal{D}^{SRP}_{ind}\right|=4N_{1}$ additional independent approaching trajectories into $\mathcal{D}^{F}$ to form the mixture dataset $\mathcal{D}^{Mix}$ according to Eq. 4 during model training, to maximize the contribution of the expensive full-stage data.
Figure 7: Model performance with the proportion of SRP data. All models are trained with the same amount of total trajectories.
4.3.3 Data Balancing Strategies
During model training, we observed that repeating the full-stage trajectories can lead to a significant better performance. To verify which repeating method is best, we compared different ways of data repeating methods in Tab. 4. In all models the announced trajectories are duplicated $\left\lfloor\frac{N_{2}}{N_{1}}\right\rfloor$ times. The results suggest that repeating the PIP segments of the full-stage data $\tau^{PIP}\subset\tau^{F}$ excluding the SRP segments, yields optimal results. This approach maintains distribution consistency among sub-tasks and implicitly regularizes the adaptation of sub-task weights, preserving the temporal dependencies between sub-tasks and stabilizing the model optimization process.
Test Scenes Zero-Shot
No Repeating 0.58 0.46
Repeat $\tau^{F}$ 0.59 0.41
Repeat $\tau^{PIP}$ 0.59 0.58
Table 4: The performance of of models trained with different PIP segment repeating methods. All models use $66\%$ of PIP segments during training.
5 Conclusion
We propose DataPlatter, a stage-decoupled training paradigm that enhances VLA models through strategic utilization of additional cost-effective spatial reasoning data. Our key contributions are threefold: (1) Additional SRP data introduced to the decoupled costly full-stage trajectories acts as a catalytic role that can achieve a 41% improvement in zero-shot success rate, by allowing enhanced training on spatial search patterns. This result, from another perspective, reduced the dependency on human-collected data. (2) The SRP/PIP mixture ratio follows a logarithmic law indicating at most 4x more additional SRP data can be added into the full-stage data to maximize the generalization performance. (3) The half-trajectories with only SRP data can serve as the bridge to transfer the generalized skills to novel target objects, providing a novel way in instance-wide generalization. Our stage-decoupled paradigm opens new directions for stage-aware curriculum learning in embodied AI, particularly in adaptive stage boundary detection.
The limitations of this work lies in (1) rigid stage segmentation struggles with multi-phase tasks which limiting its range of application, and (2) cross-object skill transfer remains suboptimal. Future work could focus on enhancing generalization on across various task and instances generalization utilizing meta-knowledge combination derived from low-cost data.
Appendix A Training Data
A.1 Our Dataset
Our dataset are collected in the simulation environment of IsaacSim, which is automatically collected using algorithm similar with Zheng et al. [50]. Each scene is initialized with 4 to 6 objects placed randomly on a table, both in position and orientation. A Franka-Panda 7-DoF robotic arm equipped with a two-finger gripper is initialized with a random end-effector pose. A static camera positioned in front of the table, along with a wrist camera mounted on the gripper, are used to capture the RGB and depth observation of the scene, as shown in Fig. 1(b). During the collection process, a target is sampled from the objects on the table and assigned as the target, and a language instruction is generated using pre-defined templates. At each step, the pose of the gripper, action targets generated by the algorithm, robot joint information, gripper status, images from the cameras, task instruction, and status information of all objects in the scene, are recorded for training and reproducing. We use ray-tracing renderer while generating camera images and evaluation. The target objects used in our experiments are shown in  Fig. 8.
Figure 8: Target objects used in our experiments.
For full-stage trajectories, we first sample collision-free grasp labels on the target, which is densely labeled using the collision model of the object using method of Fang et al. [7]. The agent then performs 6-DoF path planning using CuRobo [39] and executes the generated path. For trajectories that only involve SRP stage, to provide a implementable pipeline in real-world robots, we did not directly read the target information from the simulator. Following the method provided in Fig. 1(c), we first located the target from the RGB image captured by the static camera, after which it is feed to a CNN to detect the target bounding box. With the bounding box, we can acquire the mean depth of the target from the depth image, and calculate its position using the intrinsic and extrinsic of the camera. Then in the approach pose sampling stage, we simply samples an end-effector pose within a range of $10cm$ from the target position, ensuring the gripper is oriented towards the target, after which the paths are planned with spatial occupancy information provided by the depth image, and finally the path is executed by the agent.
A.2 CALVIN Dataset
In our datasets we use 5 tasks in D-D split of CALVIN dataset, and the tasks are shown in Tab. 5. In the dataset, each trajectory is divided into SRP and PIP segments using the methods outlined in Eq. 3. In the experiments presented in Tab. 1, we use a total of 150 trajectories per task, setting the parameter $p_{SRP}=50\%$ for the model trained with our method.
# Task
1 lift_blue_block_table
2 lift_red_block_table
3 lift_pink_block_table
4 move_slider_left
5 move_slider_right
Table 5: Tasks we used in the CALVIN dataset.
Appendix B Experiment Details
B.1 Additional Experiments
$N_{2}$ Seen Targets Novel Targets
50 0.235 0.2
100 0.375 0.275
200 0.3467 0.3025
Table 6: Generalization performance of models trained with additional data of novel targets. We set $N_{1}=100,N_{2}=0$ for seen targets and $N_{1}=50$ for novel targets.
To assess the feasibility of integrating data on novel target objects into existing comprehensive datasets collected with substantial effort, we trained models using a dataset consisting of 10 novel objects, to serve as a supplementary experiment to the results shown in Tab. 2. This dataset included a limited amount of full-stage data and a large proportion of SRP data, complemented by another 10 target objects with complete full-stage data. The results in zero-shot scenes are shown in Tab. 6, and the detailed result is shown in Tab. 7. A similar conclusion can be drawn as demonstrated in Fig. 3: by incorporating more SRP data along with a small proportion of full-stage trajectories, we can enhance performance on novel targets.
Target Object $N_{2}=50$ $N_{2}=100$ $N_{2}=200$
TS ZS TS ZS TS ZS
Seen Targets ($N_{1}=100,N_{2}=0$) Apple 0.6 0.2 0.75 0.45 0.7 0.5
Banana 0.75 0.2 0.8 0.35 0.8 0.15
Bottle 0.75 0.15 0.8 0.45 0.85 0.3
Can 0.8 0.4 0.8 0.35 0.65 0.35
Toothpaste Can 0.55 0.25 0.6 0.35 0.65 0.2
Component 0.655 0.4 0.75 0.5 0.6 0.4167
Hippo 0.45 0.15 0.6 0.35 0.5 0.5
Lock 0.75 0.25 0.75 0.65 0.75 0.55
Soap Box 0.4 0.1 0.4 0.1 0.2 0.25
White Mouse 0.35 0.25 0.45 0.2 0.6 0.25
Average 0.6055 0.235 0.67 0.375 0.63 0.3467
Novel Targets ($N_{1}=50$) Cleansing Foam 0.65 0 0.35 0.15 0.25 0.25
Hand Cream 0.4 0.15 0.6 0.2 0.4 0.2
Lemon 0.65 0.35 0.7 0.45 0.85 0.45
Mouth Rinse 0.5 0.2 0.75 0.25 0.55 0.3
Shampoo (1) 0.6 0.2 0.55 0.25 0.7 0.225
Shampoo (2) 0.65 0.2 0.55 0.3 0.5 0.35
Shampoo (3) 0.5 0.2 0.65 0.3 0.65 0.325
Shampoo (4) 0.5 0.2 0.6 0.15 0.45 0.2
Shampoo (5) 0.55 0.3 0.7 0.4 0.5 0.475
Soap 0.55 0.2 0.5 0.3 0.4 0.25
Average 0.555 0.2 0.595 0.275 0.525 0.3025
Table 7: Detailed results of models trained with additional data of novel targets.
B.2 Detailed Experiment Results
The success rate of each task in each model presented in Sec. 4 are shown from Tab. 8 to Tab. 14.
Target Object $N_{1}=50$ $N_{1}=100$ $N_{1}=200$ $N_{1}=300$ $N_{1}=400$ $N_{1}=500$
TS ZS TS ZS TS ZS TS ZS TS ZS TS ZS
Plum 0.5 0.4 0.35 0.75 0.75 0.7 0.8 0.9 0.8 0.9 0.95 0.9
Lemon 0.5 0.15 0.55 0.05 0.65 0.25 0.75 0.55 0.846 0.875 0.65 0.85
Orange 0.35 0 0.25 0.15 0.85 0.25 0.9 0.4 0.9 0.6 0.9 0.7
Apple 0.425 0.1 0.25 0.2 0.75 0.5 0.8 0.75 0.95 0.8 0.85 0.75
Peach 0.4 0.15 0.45 0.25 0.55 0.45 0.65 0.8 0.8 0.85 0.8 0.75
Average 0.435 0.16 0.37 0.28 0.71 0.43 0.78 0.68 0.8592 0.805 0.83 0.79
Table 8: Detailed results of upper bound model performance used in Fig. 3. (5 Tasks, $N_{2}=0$)
Target Object $N_{1}=100$ $N_{1}=200$ $N_{1}=300$ $N_{1}=400$
TS ZS TS ZS TS ZS TS ZS
Apple 0.7 0.35 0.8 0.5 0.85 0.55 0.85 0.55
Banana 0.8 0.35 0.7 0.4 0.65 0.45 0.8 0.85
Bottle 0.9 0.3 0.7 0.45 0.8 0.5 0.7179 0.6
Can 0.6 0.3 0.8 0.65 0.8 0.55 0.7 0.6
Toothpaste Can 0.7 0.35 0.65 0.4 0.65 0.4 0.6 0.45
Cleansing Foam 0.8 0.15 0.55 0.15 0.75 0.4 0.85 0.6
Component 0.864 0.35 0.65 0.4 0.75 0.65 0.75 0.45
Hand Cream 0.55 0.2 0.5 0.45 0.7 0.4 0.8 0.4
Hippo 0.6 0.275 0.7 0.35 0.75 0.55 0.65 0.75
Lemon 0.7 0.4 0.9 0.6 0.95 0.6 0.9 0.85
Lock 0.65 0.3 0.75 0.4 0.9 0.7 0.7 0.8
Mouth Rinse 0.8 0.4 0.6 0.5 0.9 0.5 0.9 0.5
Shampoo (1) 0.8 0.25 0.55 0.4 0.7 0.5 0.75 0.55
Shampoo (2) 0.711 0.1 0.7037 0.4 0.6 0.4554 0.6 0.55
Shampoo (3) 0.6 0.3 0.7 0.425 0.65 0.45 0.85 0.375
Shampoo (4) 0.7 0.2 0.7 0.425 0.8 0.5 0.65 0.55
Shampoo (5) 0.75 0.35 0.65 0.7 0.55 0.65 0.6 0.6
Soap 0.85 0.3 0.9 0.6 0.8 0.55 0.8696 0.4
Soap Box 0.25 0.2 0.4872 0.25 0.45 0.35 0.5 0.45
White Mouse 0.5 0.25 0.65 0.3 0.7 0.45 0.75 0.35
Average 0.6913 0.2838 0.682 0.4375 0.735 0.5128 0.7394 0.5413
Table 9: Detailed results of upper bound model performance used in Fig. 3. (20 Tasks, $N_{2}=0$)
Target Object $N_{2}=100$ $N_{2}=200$ $N_{2}=300$ $N_{2}=400$
TS ZS TS ZS TS ZS TS ZS
5 Tasks Plum 0.65 0.55 0.65 0.8 0.6 0.75 0.7 0.75
Lemon 0.6 0.1 0.65 0.55 0.6625 0.85 0.685 0.685
Orange 0.55 0.5 0.7 0.35 0.6 0.55 0.65 0.55
Apple 0.5 0.2 0.65 0.5 0.7 0.6 0.85 0.7
Peach 0.5 0.5 0.3 0.7 0.67 0.6 0.575 0.7
Average 0.56 0.37 0.59 0.58 0.6925 0.67 0.712 0.695
20 Tasks Apple 0.55 0.2 0.55 0.5 0.4 0.4 - -
Banana 0.75 0.2 0.75 0.325 0.65 0.55 - -
Bottle 0.65 0.25 0.775 0.55 0.7 0.3 - -
Can 0.6 0.15 0.6 0.4667 0.5 0.45 - -
Toothpaste Can 0.6 0.2 0.5 0.35 0.5 0.15 - -
Cleansing Foam 0.5 0.25 0.65 0.1 0.5 0.6 - -
Component 0.65 0.1 0.7 0.45 0.65 0.25 - -
Hand Cream 0.5 0.45 0.75 0.2 0.45 0.25 - -
Hippo 0.65 0.35 0.55 0.5 0.55 0.75 - -
Lemon 0.65 0.4 0.95 0.55 0.75 0.75 - -
Lock 0.35 0.35 0.85 0.7 0.45 0.45 - -
Mouth Rinse 0.65 0.3 0.55 0.45 0.75 0.35 - -
Shampoo (1) 0.6 0.2 0.65 0.45 0.55 0.4 - -
Shampoo (2) 0.7 0.35 0.7 0.3448 0.7 0.45 - -
Shampoo (3) 0.55 0.4 0.7027 0.35 0.7 0.5 - -
Shampoo (4) 0.55 0.15 0.6 0.5 0.5 0.35 - -
Shampoo (5) 0.7 0.55 0.65 0.55 0.65 0.65 - -
Soap 0.5 0.3 0.65 0.4 0.6 0.45 - -
Soap Box 0.25 0.15 0.4 0.3 0.6 0.3 - -
White Mouse 0.4 0.2 0.65 0.25 0.35 0.25 - -
Average 0.7675 0.275 0.6589 0.4143 0.575 0.43 - -
Table 10: Detailed results of models trained with different amount of SRP data used in Fig. 3. ($N_{1}=100$)
Target Object w/o SRP ($N_{1}=100,N_{2}=0$) w/ SRP ($N_{1}=100,N_{2}=200$) UB ($N_{1}=300,N_{2}=0$)
TS ZS TS ZS TS ZS
Plum 0.3 0.25 0.35 0.35 0.85 0.7
Lemon 0.1364 0.05 0.45 0.35 0.6 0.35
Orange 0.15 0.05 0.3 0.1 0.3 0.2
Apple 0.15 0.1 0.35 0.1 0.55 0.4
Peach 0.2 0.2 0.35 0.5 0.5 0.55
Average 0.1873 0.13 0.36 0.28 0.56 0.44
Table 11: Detailed results of Tab. 1. Performance of RoboFlamingo [22] on our dataset.
Target Object w/o SRP ($N_{1}=75,N_{2}=0$) w/ SRP ($N_{1}=75,N_{2}=75$) UB ($N_{1}=150,N_{2}=0$)
RM RF RM RF RM RF
lift_blue_block_table 0.4583 0.3333 0.5833 0.2917 0.7717 0.5
lift_pink_block_table 0.6 0.16 0.92 0.48 0.88 0.6
lift_red_block_table 0.3333 0.0833 0.625 0.2083 0.9183 0.1667
move_slider_left 0.9219 1 0.9219 1 0.9244 1
move_slider_right 1 1 1 1 1 1
Average 0.8036 0.7366 0.8839 0.7812 0.9254 0.8125
Table 12: Detailed results of Tab. 1. Performance of RoboMM (RM) [47] and RoboFlamingo (RF) [22] on CALVIN.
Object Group Target Object None ($N_{1}=0,N_{2}=0$) Only SRP ($N_{1}=0,N_{2}=200$) Full-Stage ($N_{1}=200,N_{2}=0$)
Seen $N_{1}=200$ $N_{2}=0$ Apple 0.65 0.5 0.5
Lemon 0.65 0.7 0.6
Orange 0.75 0.7 0.65
Peach 0.65 0.7 0.8
Average 0.675 0.65 0.6375
Novel Plum 0.05 0.4 0.65
Table 13: Detailed results of Tab. 2. Performance of the models on seen and novel target objects in test scenes. (5 Tasks)
Object Group Target Object None ($N_{1}=0,N_{2}=0$) Only SRP ($N_{1}=0,N_{2}=100$) Full-Stage ($N_{1}=100,N_{2}=0$)
Seen $N_{1}=100$ $N_{2}=0$ Apple 0.35 0.4 0.7
Banana 0.6 0.8 0.8
Bottle 0.65 0.825 0.9
Can 0.7 0.65 0.6
Toothpaste Can 0.65 0.65 0.7
Cleansing Foam 0.6 0.55 0.8
Component 0.7 0.5 0.864
Hippo 0.5 0.5 0.6
Shampoo (2) 0.7 0.7 0.711
Shampoo (4) 0.6 0.5 0.7
Average 0.605 0.6075 0.7375
Novel Hand Cream 0 0.2 0.55
lemon 0.05 0.225 0.7
Lock 0.05 0.15 0.65
Mouth Rinse 0 0.3 0.8
Shampoo (1) 0 0.15 0.8
Shampoo (3) 0.05 0.175 0.6
Shampoo (5) 0.05 0.45 0.75
Soap 0 0.075 0.85
Soap Box 0 0.15 0.25
White Mouse 0.05 0.15 0.5
Average 0.025 0.2025 0.645
Table 14: Detailed results of Tab. 2. Performance of the models on seen and novel target objects in test scenes. (20 Tasks)